rb of 1 mw al 93 18 2 127 211 eal pie 20, 20, 77 H 0 0 2 212 119 2 QQ ot M e rd xad 9 m pp to pptt
Answer #1
Sr, No X (X - X̄)² sorted X
1 18.71 4.233 18.04
2 21.41 0.413 18.71
3 20.72 0.002 18.92
4 21.81 1.087 19.25
5 19.29 2.183 19.29
6 22.43 2.764 19.44
7 20.17 0.357 19.77
8 23.71 8.658 20.17
9 19.44 1.762 20.33
10 20.5 0.072 20.5
11 18.92 3.413 20.72
12 20.33 0.191 21.12
13 23 4.984 21.41
14 22.85 4.337 21.77
15 19.25 2.303 21.81
16 21.77 1.005 22.11
17 22.11 1.802 22.43
18 19.77 0.995 22.85
19 18.04 7.439 23
20 21.12 0.124 23.71
X (X - X̄)²
total sum 415.35 48.125375
n 20 20

a)

mean =    ΣX/n =    20.7675
b)

Median = (n+1)/2 = 10.5th value of sorted data

so, median = 20.61

c)

10% trimmed mean

20 numbers are there

10% 0f 20 = 2

so removing the two lowest and two highest numbers and take their average

18.92
19.25
19.29
19.44
19.77
20.17
20.33
20.5
20.72
21.12
21.41
21.77
21.81
22.11
22.43
22.85

10% trimmed mean= 20.7431

d)

Q1=1/4(n+1)=21/4=5.25th value = 19.29+(19.29-19.44)*0.25=19.3275

Q3=0.75(n+1)=15.75th value = 21.81+0.75*(22.11-21.81)=22.035

e)

IQR = Q3-Q1 = 2.7075

f)

lower bound=Q1-1.5IQR = 15.26625

g) upper bound = Q3+1.5*IQR = 26.0962

h)

25 23.71 22.035 20.61 19.3275 20.7675 20 18.04 15

i)

maximum=23.71

j) minimum=18.04

k) Range=5.67

l) sample variance =    Σ(X - X̄)²/(n-1)=   2.533

m) sample std dev =   √ [ Σ(X - X̄)²/(n-1)] =   1.5915